# Mixing Artificial and Natural Intelligence: From Statistical Mechanics to AI and Back to Turbulence

### v1

@@ 4.8 Applied Mathematics (and Theoretical/Statistical Physics) for AI: What is Next?
  • State-Space Structures in AI Diffusion Models: Harnessing PDEs and Graph Theory for Advanced Capabilities: The diffusion models of AI, often employed in the rendering of two-dimensional imagery (and potentially high-dimensional snapshots/images), have largely not exploited the intricate structure of state-space. Such structures, implicit in the juxtaposition of image pixels or the manifold of higher-dimensional data, beckon for a more sophisticated incorporation into the AI models. It is reasonable to prognosticate that the field would undergo a significant transformation with the systematic integration of these state-space structures, particularly through the lens of Partial Differential Equations (PDEs) that embrace both spatial and temporal dimensions (d+1), as well as through the application of graph theory and graphical models Wainwright & Jordan (2007); Chertkov (2024), expressing relations (deterministic and stochastic) of the underlying applications. These mathematical structures, when finely tuned and interwoven with diffusion processes, hold the potential to significantly expand the capabilities of AI diffusion models, advancing them into previously uncharted and complex application realms.
@@ 5.1 Scalar Turbulence: From Fields (Eulerian) to Particles (Lagrangian)
The advection of the passive scalar field, $\theta(t;{\bm{r}})$, which evolves in space ${\bm{r}}\in\mathbb{R}^{d}$ and time $t$, is governed by the following Partial Differential Equation (PDE) Kraichnan (1968); Batchelor (1959)
@@ 5.2 Tetrad Model: Lagrangian Phenomenology
where $\bm{W}_{M}(t)$ and $\bm{W}_{g}(t)$ are zero-mean stochastic terms modeled as Wiener processes, with their covariances dependent on $\bm{M}$ and $\bm{g}$. The covariance matrices and the tuning parameter $\alpha$ in Eqs. (14,15) were empirically calibrated using post-factum Eulerian Direct Numerical Simulations (DNS), as detailed in Chertkov et al. (1999). Further calibration of the tetrad phenomenology against Lagrangian data was performed in Pumir et al. (2000).
At this stage in the story it seems appropriate to make a step back and describe the rich pre-history which has led to development of the tetrad model (14,15). Indeed, exploration of the Velocity Gradient Tensor (VGT) within the realm of developed incompressible turbulence has been a subject of enduring interest and investigation since the early 1980s. The early closure attempts of the NS Eq. (12), first by Vieillefosse (1982) and then by Cantwell (1992) led to the Restricted Euler (RE) equation – which is just Eq. (13) with zero right hand side. However, this, so-called Viellefosse, equation was clearly problematic due to its nonphysical finite time singularity. Concurrently, and leading up to Chertkov et al. (1999), Girimaji & Pope (1990) embarked on exploring related stochastic models and phenomenology for the VGT. These efforts, focused on the local dynamics of the VGT (rather than its coarse-grained behavior which came later in Chertkov et al. (1999)) introduced stochastic modifications to the right-hand side of Eq. (13). The stochastic terms were interpreted as immitating effect of larger scales on smaller scales.
@@ 5.3 Pressure Hessian as a Function of Velocity Gradient: Lagrangian Neural Closure
Tian et al. (2021) suggested to represent ${\bm{H}}$ as a function of ${\bm{m}}$ via a Deep Neural Networks (DNNs) trained on the data from the Direct Numerical Simulations (DNS) of turbulence. This is why we call the approach – Lagrangian Neural Closure (LNC) for the pressure Hessian. Moreover, and utilizing earlier ideas of Pope (1975), also developed by Ling et al. (2016)), the DNN-parameterized function (see Fig. (6)) was constrained to obey rotational symmetry of the homogeneous isotropic turbulence symmetries, then resulting, according to the Cayley-Hamilton theorem, in the finite representation via the tensor basis:
@@
In summary, Tian et al. (2021)) anchored the deterministic segment of the phenomenological model in the local approximation closures while incorporating core physical constraints – Galilean invariance (enforced automatically in the Lgrangian frame), rotational invariance, and the incompressibility condition. These principles were explicitly integrated into Eq. (13), with modifications such as replacing ${\bm{h}}$ according to Eq. (17), and analogous alterations for injection and dissipation terms. The model was further enhanced by a stochastic term with zero mean, representing the impact of scale fluctuations on the principal viscous scale interactions. The approach harnessed Deep Neural Networks (DNNs) to capture functional degrees of freedom, like $g^{(n)}(\cdot)$ from Eq. (17), through training on Lagrangian data derived from high-Reynolds number DNS. The validation process involved a rigorous out-of-sample test, which demonstrated the model’s enhanced ability to represent the magnitude and orientation of the small-scale pressure Hessian contributions. The joint probability density function of the second and third invariants of the Velocity Gradient Tensor was in a satisfactory concordance with the DNS ’ground-truth’ data. Moreover, the model successfully captured several key turbulence structures, even those not explicitly embedded within it.
@@ 5.4 Evolution of AI for Turbulence (Eulerian and also Lagrangian)
In the latest advancements of the Generative Turbulence (GenTurb) methodology, AI diffusion models have been employed to supplement, enhance, or generate samples of turbulent flows in both Eulerian and Lagrangian frameworks, as detailed in Li et al. (2023a) and Li et al. (2023b) respectively.
@@
This subsection concludes with two examples that diverge from the previously discussed categories of Neural Closure, GenTurb, and Neural Turb-ODE, yet exhibit some similarities. Fonda et al. (2019) suggest using DNNs to compress flow snapshots indicative of convection, utilizing a U-shaped DNN architecture with contraction and expansion branches to distill the complex 3D turbulent structure into a simpler temporal planar representation at the convection layer’s midplane. Stulov & Chertkov (2021) develop a Neural Particle Image Velocimetry (PIV) technique to reconstruct the velocity field from consecutive snapshots of the density (passive scalar) field. While acknowledging the field’s breadth and the emergence of recent works like, Han, Yuhuan & Wang, Quian (2023) and Yu et al. (2023) which apply transformers to similar ends, we note a commonality between the problems of snapshot compression and velocity reconstruction from scalar field snapshots. Both entail learning an instantaneous mapping between two related but distinct snapshots of physical fields. This shared characteristic suggests grouping these instances (and potentially others) into a category termed Neural Maps.
@@ 5.5 From Smooth Particle Hydrodynamics to Neural Lagrangian Large Eddy Simulation
The model presented in Tian et al. (2023) was trained using Ground Truth (GT) data derived from Direct Numerical Simulations (DNS) on a $1,024^{3}$ Eulerian grid, alongside Lagrangian trajectories of $64^{3}$ non-inertial fluid particles carried by the DNS flow. The dataset employed in Woodward et al. (2023), while smaller in scale, was fundamentally similar in its approach and structure.
Figure 8: Performance on the trained L-LES model in interpolation and extrapolation errors over delayed intervals and different level of turbulent Mach number, respectively. (a) shows normalized $L_{2}$ error; and (b) shows Probability Distribution Functions (PDF) of the prediction.
@@
In contrast, Woodward et al. (2023) conducted the training of the Neural Lagrangian models in a single stage, utilizing a loss function that integrated two main components: (a) matching the velocity field across the Eulerian grid and (b) KL matching of single-particle Lagrangian velocity distributions. Additionally, Woodward et al. (2023) developed a mixed-mode method that combines local sensitivity analysis with forward and backward automatic differentiation to efficiently compute gradients of the loss function across both physical and DNN-embedded parameters.
@@ 5.6 AI for Turbulence: What is Next?
  • Less Data – More Physics: I suggest that the integration of physics-guided principles with AI diffusion models can significantly enhance data compression in multi-scale scientific challenges, thereby improving efficiency and accuracy. Leveraging the advancements in AI diffusion, transformers, and reinforcement learning (RL) discussed in this paper, along with other cutting-edge AI methods, we aim to refine these models for optimal data compression. This enhancement should maintain or even augment the integrity and interpretability of critical features, akin to the Lagrangian LES models previously mentioned. Furthermore, adopting an RL approach to data reduction, by assigning a "price tag" to data based on its utility for specific physical phenomena or tasks in turbulence research, could prioritize data in a way that enhances data acquisition efficiency. Integrating Uncertainty Quantification (UQ) methods should further strengthen performance evaluation and support more strategic decision-making in both the discovery and application aspects of turbulence studies.
  • Diffusion Models of AI with Physical Time Line: Aligning with Holzschuh et al. (2023), adapting AI diffusion models to physical time processes appears highly promising, transitioning from artificial to physical time dynamics. In this context, expanding the Lagrangian perspective on turbulence, as detailed in previous sections, to encompass fluid blob dynamics that transport correlations from large integral scales to smaller scales of interest in the inertial range, is a viable path. This approach would facilitate the integration of AI into turbulence research, particularly for analyzing Eulerian field statistics like velocity or scalar fields.

### v2

@@ 4.8 Applied Mathematics (and Theoretical/Statistical Physics) for AI: What is Next?
  • State-Space Structures in Diffusion Models: Harnessing PDEs and Graph Theory for Advanced Capabilities: The diffusion models of AI, often employed in the rendering of two-dimensional imagery (and potentially high-dimensional snapshots/images), have largely not exploited the intricate structure of state-space. Such structures, implicit in the juxtaposition of image pixels or the manifold of higher-dimensional data, beckon for a more sophisticated incorporation into the AI models. It is reasonable to prognosticate that the field would undergo a significant transformation with the systematic integration of these state-space structures, particularly through the lens of PDEs that embrace both spatial and temporal dimensions (d+1), as well as through the application of graph theory and graphical models Wainwright & Jordan (2007); Chertkov (2024), expressing relations (deterministic and stochastic) of the underlying applications. These mathematical structures, when finely tuned and interwoven with diffusion processes, hold the potential to significantly expand the capabilities of diffusion models, advancing them into previously uncharted and complex application realms.
@@ 5.1 Scalar Turbulence: From Fields (Eulerian) to Particles (Lagrangian)
The advection of the passive scalar field, $\theta(t;{\bm{r}})$, which evolves in space ${\bm{r}}\in\mathbb{R}^{d}$ and time $t$, is governed by the following PDE Kraichnan (1968); Batchelor (1959)
@@ 5.2 Tetrad Model: Lagrangian Phenomenology
where $\bm{W}_{M}(t)$ and $\bm{W}_{g}(t)$ are zero-mean stochastic terms modeled as Wiener processes, with their covariances dependent on $\bm{M}$ and $\bm{g}$. The covariance matrices and the tuning parameter $\alpha$ in Eqs. (14,15) were empirically calibrated using post-factum Eulerian DNS, as detailed in Chertkov et al. (1999). Further calibration of the tetrad phenomenology against Lagrangian data was performed in Pumir et al. (2000).
At this stage in the story it seems appropriate to make a step back and describe the rich pre-history which has led to development of the tetrad model (14,15). Indeed, exploration of the Velocity Gradient Tensor (VGT) within the realm of developed incompressible turbulence has been a subject of enduring interest and investigation since the early 1980s. The early closure attempts of the NS Eq. (12), first by Vieillefosse (1982) and then by Cantwell (1992) led to the Restricted Euler (RE) equation – which is just Eq. (13) with zero right hand side. However, this, so-called Vieillefosse equation was clearly problematic due to its nonphysical finite time singularity. Concurrently, and leading up to Chertkov et al. (1999), Girimaji & Pope (1990) embarked on exploring related stochastic models and phenomenology for the VGT. These efforts, focused on the local dynamics of the VGT (rather than its coarse-grained behavior which came later in Chertkov et al. (1999)) introduced stochastic modifications to the right-hand side of Eq. (13). The stochastic terms were interpreted as immitating effect of larger scales on smaller scales.
@@ 5.3 Pressure Hessian as a Function of Velocity Gradient: Lagrangian Neural Closure
Tian et al. (2021) suggested to represent ${\bm{H}}$ as a function of ${\bm{m}}$ via a DNNs trained on the data from the Direct Numerical Simulations (DNS) of turbulence. This is why we call the approach – Lagrangian Neural Closure (LNC) for the pressure Hessian. Moreover, and utilizing earlier ideas of Pope (1975), also developed by Ling et al. (2016)), the DNN-parameterized function (see Fig. (6)) was constrained to obey rotational symmetry of the homogeneous isotropic turbulence symmetries, then resulting, according to the Cayley-Hamilton theorem, in the finite representation via the tensor basis:
@@
In summary, Tian et al. (2021)) anchored the deterministic segment of the phenomenological model in the local approximation closures while incorporating core physical constraints – Galilean invariance (enforced automatically in the Lgrangian frame), rotational invariance, and the incompressibility condition. These principles were explicitly integrated into Eq. (13), with modifications such as replacing ${\bm{h}}$ according to Eq. (17), and analogous alterations for injection and dissipation terms. The model was further enhanced by a stochastic term with zero mean, representing the impact of scale fluctuations on the principal viscous scale interactions. The approach harnessed DNNs to capture functional degrees of freedom, like $g^{(n)}(\cdot)$ from Eq. (17), through training on Lagrangian data derived from high-Reynolds number DNS. The validation process involved a rigorous out-of-sample test, which demonstrated the model’s enhanced ability to represent the magnitude and orientation of the small-scale pressure Hessian contributions. The joint probability density function of the second and third invariants of the Velocity Gradient Tensor was in a satisfactory concordance with the DNS ’ground-truth’ data. Moreover, the model successfully captured several key turbulence structures, even those not explicitly embedded within it.
@@ 5.4 Evolution of AI for Turbulence (Eulerian and also Lagrangian)
In the latest advancements of the Generative Turbulence (GenTurb) methodology, diffusion models have been employed to supplement, enhance, or generate samples of turbulent flows in both Eulerian and Lagrangian frameworks, as detailed in Li et al. (2023a) and Li et al. (2023b) respectively.
@@
This subsection concludes with two examples that diverge from the previously discussed categories of Neural Closure, GenTurb, and Neural Turb-ODE, yet exhibit some similarities. Fonda et al. (2019) suggest using DNNs to compress flow snapshots indicative of convection, utilizing a U-shaped DNN architecture with contraction and expansion branches to distill the complex 3D turbulent structure into a simpler temporal planar representation at the convection layer’s midplane. The U-net architecture was also utilized by Cremades et al. (2024) to predict the emergence of coherent structures in boundary layer flows. Stulov & Chertkov (2021) develop a Neural Particle Image Velocimetry (PIV) technique to reconstruct the velocity field from consecutive snapshots of the density (passive scalar) field. While acknowledging the field’s breadth and the emergence of recent works like, Han, Yuhuan & Wang, Quian (2023) and Yu et al. (2023) which apply transformers to similar ends, we note a commonality between the problems of snapshot compression and velocity reconstruction from scalar field snapshots. Both entail learning an instantaneous mapping between two related but distinct snapshots of physical fields. This shared characteristic suggests grouping these instances (and potentially others) into a category termed Neural Maps.
@@ 5.5 From Smooth Particle Hydrodynamics to Neural Lagrangian Large Eddy Simulation
The model presented in Tian et al. (2023) was trained using Ground Truth (GT) data derived from DNS on a $1,024^{3}$ Eulerian grid, alongside Lagrangian trajectories of $64^{3}$ non-inertial fluid particles carried by the DNS flow. The dataset employed in Woodward et al. (2023), while smaller in scale, was fundamentally similar in its approach and structure.
Figure 8: (From Tian et al. (2023).) Performance on the trained L-LES model in interpolation and extrapolation errors over delayed intervals and different level of turbulent Mach number, respectively. (a) shows normalized $L_{2}$ error; and (b) shows Probability Distribution Functions (PDF) of the prediction.
@@
In contrast, Woodward et al. (2023) trained the Neural Lagrangian models in a single stage using a loss function that integrated two main components: (a) matching the velocity field across the Eulerian grid and (b) KL matching of single-particle Lagrangian velocity distributions. Additionally, Woodward et al. (2023) developed a mixed-mode method combining local sensitivity analysis with forward and backward AD to efficiently compute gradients of the loss function across both physical and DNN-embedded parameters. 555 AD can be performed in two primary modes: forward and reverse. Forward mode AD is efficient for functions with fewer inputs and more outputs, propagating derivatives alongside the original function evaluation. In contrast, reverse mode AD, also known as backpropagation, is optimal for functions with many inputs and a single output, making it particularly useful for DNNs. In the context of the Neural Lagrangian models, reverse mode AD facilitates the optimization of the loss function by efficiently computing gradients with respect to both physical and DNN-embedded parameters. For further insight on modern ways of utilizing AD in learning-based, simulation-driven scientific discovery, refer to Lavin et al. (2022).
@@ 5.6 AI for Turbulence: What is Next?
  • Less Data – More Physics: I suggest that the integration of physics-guided principles with diffusion models can significantly enhance data compression in multi-scale scientific challenges, thereby improving efficiency and accuracy. Leveraging the advancements in diffusion, transformers, and reinforcement learning (RL) discussed in this paper, along with other cutting-edge AI methods, we aim to refine these models for optimal data compression. This enhancement should maintain or even augment the integrity and interpretability of critical features, akin to the Lagrangian LES models previously mentioned. Furthermore, adopting an RL approach to data reduction, by assigning a "price tag" to data based on its utility for specific physical phenomena or tasks in turbulence research, could prioritize data in a way that enhances data acquisition efficiency. Integrating Uncertainty Quantification (UQ) methods should further strengthen performance evaluation and support more strategic decision-making in both the discovery and application aspects of turbulence studies.
  • Diffusion Models of AI with Physical Time Line: Aligning with Holzschuh et al. (2023), adapting diffusion models to physical time processes appears highly promising, transitioning from artificial to physical time dynamics. In this context, expanding the Lagrangian perspective on turbulence, as detailed in previous sections, to encompass fluid blob dynamics that transport correlations from large integral scales to smaller scales of interest in the inertial range, is a viable path. Notice that most recently an interesting step in this direction was taken in Li et al. (2024a; b), applying diffusion models to the Lagrangian evolution of particles in turbulent flows. A machine learning diffusion model that generates single-particle trajectories was developed. Even though this diffusion model was not physics-informed, it seems to reproduce statistical benchmarks of Lagrangian turbulence and even demonstrate surprisingly strong generalizability for extreme events. Further development of this approach and its integration with physics-informed modeling, including Eulerian field statistics like velocity or scalar fields, should further facilitate the integration of the diffusion modeling into turbulence research.
  • Transformer Models in Turbulence Modeling: Transformers, as we briefly discussed in Section 2.4, represent a branch of generative modeling that has led to some of the most remarkable recent advancements in AI, particularly in the realm of Large Language Models (LLM). Zhu et al. (2024) introduced Fluid-LLM, which tokenizes a snapshot and then a sequence of snapshots of the fluid flow, using it as a query to generate predictions for how the flow evolves. Solera-Rico et al. (2023) applied $\beta$-Variational Autoencoders (which can be viewed as early versions of transformers) and transformers themselves for reduced-order modeling of fluid flows, highlighting these models’ capability to efficiently capture complex flow dynamics with reduced computational costs. Yousif et al. (2023) developed a transformer-based synthetic-inflow generator specifically designed for spatially developing turbulent boundary layers. Their work demonstrated the effectiveness of transformers in generating high-fidelity inflow conditions, which are crucial for accurate turbulence simulations. Notably, these applications of transformer-related technology are Eulerian by construction. We envision extending transformer technology to the Lagrangian frameworks, such as the Neural Lagrangian Large Eddy Simulation discussed in Section 5.5, by using Lagrangian quasi-particles as tokens.
@@
  • New Hypotheses: Turbulence is often referred to as the last unsolved problem of classical mechanics. Over the multi-century history of the field, many hypotheses have been proposed. While the majority have been proven wrong, quite a number were abandoned simply because we lacked the necessary validation tools. With new AI tools at our disposal, we should re-examine these hypotheses. Some of our discussions above, notably in Section 5, can be viewed as first steps in this direction. Another interesting example of this type involves developing a famous hypothesis by Hopf made in Hopf (1948), suggesting that turbulence can be explained in terms of the unstable simple invariant solutions populating the inertial manifold of a high-dimensional space. This idea was further explored in Page et al. (2024), where it was shown how to reconstruct PDFs of dissipation rate, production rate, and energy in developed two-dimensional turbulence from a set of unstable periodic orbits. A combination of AI techniques – AD and DNN (of the deep convolutional autoencoder type) – as well as Markov chain approaches were utilized to find hundreds of unstable periodic orbit solutions and to build the PDFs. We expect to see more approaches in this spirit, using AI to reconsider old and validate new bold hypotheses about turbulence.